OpenStudy (anonymous):

y1=5x+6, y2=x+9, y3=x, and the difference between 9 times y1, and 44 times y2 is 342 less that y3

OpenStudy (tkhunny):

So, 9(y1) - 44(y2) = y3 - 342 After some substitution, you should have a little algebra to do. I am worried about the signs of things. It y2 < y1 in the solution set? Is y3 - 342 > 0 in the solution set? You WILL have to check anything you think is an answer. There remain mysteries, here! You may need: So, 44(y2) - 9(y1) = y3 - 342 Do some exploring!